# Enumerative Geometry of Quantum Periods

### v1

@@ 5.3. Winding-$w$ open/closed correspondence
By flop invariance of the topological vertex [KM], we have,
(5.7) $Z_{W}(Q^{\ell_{i}},Q^{E})=Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})$
via the equality of Kähler parameters $Q^{E}=Q^{-C}$. Equivalently, Equation 5.7 is,
(5.8) $\begin{split}&\sum_{\lambda^{\prime}_{k+1}}Q^{|\lambda^{\prime}_{k+1}|E}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}\\ &=\sum_{\lambda_{k+1}}Q^{-|\lambda_{k+1}|C}\sum_{\lambda_{1},\ldots,\lambda_{k}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k}C_{\lambda^{T}_{j}\lambda_{j+1}\varnothing}\right)C_{\lambda_{k+1}^{T}\lambda_{1}\varnothing}\end{split}$
Restricting to $|\lambda^{\prime}_{k+1}|=w$, the coefficient of $Q^{wE}$ must equal the coefficient of $Q^{-wC}$. Therefore, we have the first statement in Theorem 5.3,
$\sum_{|\lambda^{\prime}_{k+1}|=w}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}=Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})|_{\pi^{*}\beta-wC}$
where $Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{C})|_{\pi^{*}\beta-wC}$ is defined in Section 5.3.
For the second statement in Theorem 5.3, we take the logarithm of the partition function to get the free energy, i.e., the generating function of connected Gromov-Witten invariants. We have,
$\log Z_{W}(Q^{\ell_{i}},Q^{E})$ $=\log\left(1+\sum_{\lambda^{\prime}_{k+1},\exists\lambda_{i}\neq 0}Q^{|\lambda^{\prime}_{k+1}|E}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}\right)$
$=\sum_{g,\beta,w\in\mathbb{Z}_{\geq 0}}N_{g}(W,\beta+wE)\hbar^{2g}Q^{\beta+wE}$
and,
$\log Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})$
$=\log\left(1+\sum_{\exists\lambda_{i}\neq 0}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}Q^{-|\lambda_{k+1}|C}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\left(\prod_{j=1}^{k}C_{\lambda^{T}_{j}\lambda_{(j+1)}\varnothing}\right)\right.\left.C_{\lambda_{k+1}^{T}\lambda_{1}\varnothing}\right)$
$=\sum_{g,\beta,w\in\mathbb{Z}_{\geq 0}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)\hbar^{2g}Q^{\beta-wC}$
The equality of partition functions $\log Z_{W}=\log Z_{\widehat{\mathbb{P}}}$ implies that the coefficient of $Q^{-|\lambda_{k+1}|C}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}$ in $\log Z_{\widehat{\mathbb{P}}}$ is equal to the coefficient of $Q^{|\lambda^{\prime}_{k+1}|E}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}$ in $\log Z_{W}$, i.e., $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)=N_{g}(W,\beta+wE)$.
We express $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ explicitly in terms of vertex functions of $Z_{W}$. We define $S:=\sum_{\begin{subarray}{c}\lambda^{\prime}_{k+1},\\ \exists\lambda_{i}\neq 0\end{subarray}}Q^{|\lambda^{\prime}_{k+1}|E}Z_{\mathbb{P}}(Q^{\ell_{i}},V)|_{\lambda^{\prime}_{k+1}}Z_{\mathbb{C}^{3}}(V)|_{\lambda^{{}^{\prime}T}_{k+1}}$, hence $\log Z_{W}=\log(1+S)=\sum_{l\geq 0}\frac{(-1)^{l+1}}{l}S^{l}$.
We find the coefficient of $Q^{\beta+wE}$ in $\log Z_{W}$. We only consider those powers $S^{l}$ where $l|w$. Let $(w_{1},\ldots,w_{l})$ be a partition of $w$ where $\sum_{i}w_{i}=w$ with $w_{i}\geq 0$. We see that the coefficient of $Q^{wE}$ in $S^{l}$ is,
$\sum_{(w_{1},\ldots,w_{l})}\sum_{\begin{subarray}{c}1\leq j\leq l,\\ |(\lambda^{\prime}_{k+1})_{j}|=w_{j}\end{subarray}}P_{1}\ldots P_{l}$
where we define $P_{1}\ldots P_{l}:=\left(\prod_{i=1}^{l}Z_{\mathbb{P}}(V)|_{(\lambda^{\prime}_{k+1})_{i}}Z_{\mathbb{C}^{3}}(V)|_{(\lambda^{{}^{\prime}T}_{k+1})_{i}}\right)$, and $(\lambda^{\prime}_{k+1})_{j}$ is the representation attached to the external leg for the $j$-th term in $P_{j}$ for $1\leq j\leq l.$
Suppose we write $\beta=c_{1}\ell_{1}+\ldots c_{k}\ell_{k}$ in the basis given by the compact 1-dimensional torus orbits. Let $(\beta_{1},\ldots,\beta_{l})$ be a partition of $\beta$, i.e., $\sum_{i=1}^{l}\beta_{i}=\beta$, and we write $\beta_{i}=c_{i1}\ell_{1}+\ldots+c_{ik}\ell_{k}$ with $\sum_{i=1}^{l}c_{ia}=c_{a}$ for $a=1,\ldots,k$, and $c_{ia}\geq 0$. For $1\leq i\leq l$, we suppose that the coefficient of $Q^{\beta_{i}}$ appears only in $P_{i}$. It is given by,
$\sum_{\begin{subarray}{c}1\leq a\leq k,\\ |\lambda_{ia}|=c_{ia}\end{subarray}}(-1)^{\sum_{a}|\lambda_{ia}|}q^{\sum_{a}\kappa_{\lambda_{ia}}}\left(\prod_{a=1}^{k-1}C_{\lambda^{T}_{ia}\lambda_{(i+1)a}\varnothing}\right)C_{\lambda^{T}_{ka}\lambda_{1a}(\lambda^{\prime}_{k+1})_{i}}C_{(\lambda^{{}^{\prime}T}_{k+1})_{i}\varnothing\varnothing}$
where $\lambda_{ia}$ for $a=1,\ldots,k$ are the partitions used in $P_{i}$. Taking the product over all $i$ gives the coefficient of $Q^{\beta}$ in $S^{l}$. Thus, also summing over $l$, the coefficient of $Q^{\beta+wE}$ is,
$\sum_{\begin{subarray}{c}l\geq 0,l|w,\\ (w_{1},\ldots,w_{l})\in\mathbb{Z}^{l}_{\geq 0},\\ 1\leq j\leq l,|(\lambda^{\prime}_{k+1})_{j}|=w_{j}\end{subarray}}\prod_{i=1}^{l}\sum_{\begin{subarray}{c}1\leq a\leq k,\\ |\lambda_{ia}|=c_{ia},\\ c_{ia}\in\mathbb{Z}_{\geq 0}\end{subarray}}\left[\frac{(-1)^{l+1+\sum_{a}|\lambda_{ia}|}}{l}q^{\sum_{a}\kappa_{\lambda_{ia}}}\left(\prod_{a=1}^{k-1}C_{\lambda^{T}_{ia}\lambda_{(i+1)a}\varnothing}\right)C_{\lambda^{T}_{ka}\lambda_{1a}(\lambda^{\prime}_{k+1})_{i}}C_{(\lambda^{{}^{\prime}T}_{k+1})_{i}\varnothing\varnothing}\right]$
Substituting $q=e^{i\hbar}$ in the above, the resulting $\hbar^{2g}$-coefficient is the invariant $N_{g}(W,\beta+wE)$, which is equal to $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ by flop invariance [KM]. ∎
Theorem 5.3 expresses the closed invariant $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)$ of local $\widehat{\mathbb{P}}$ in terms of a product of open invariants of an outer AV-brane in local $\mathbb{P}$ and open invariants of an outer AV-brane in $\mathbb{C}^{3}.$
5.4. Application: winding-1 open-closed equality
Specializing Theorem 5.3 to winding-1 open invariants, we have the following corollary equating open and closed BPS invariants in all genera.
@@ 5.4. Application: winding-1 open-closed equality
Suppose $\mathbb{P}$ is toric, and $\pi:\widehat{\mathbb{P}}\rightarrow\mathbb{P}$ is a toric blow up at a point with exceptional curve $C$. Let $K_{\mathbb{P}},K_{\widehat{\mathbb{P}}}$ be the toric canonical bundles. Then, we have
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=(-1)^{g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$
For some explicit values of open and closed BPS invariants, we refer to Section 5, [GZ] and Appendix B, [HKR] respectively.
Proof of Corollary 5.4.
We first derive a statement equivalent to Corollary 5.4. The Gopakumar-Vafa formula [GV1] [GV2] for $N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ states,
The open multiple cover formula [MV] for $O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ states,
(5.10) $\sum_{g\geq 0}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}=\sum_{g\geq 0}(-1)^{g+1}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\left(2\sin\frac{\hbar}{2}\right)^{2g-1}$
From Equation 5.10, Corollary 5.4 is equivalent to the equality,
@@
(5.11) $\sum_{g\geq 0}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}=\sum_{g\geq 0}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\left(2\sin\frac{\hbar}{2}\right)^{2g-1}$
After dividing by $\left(2\sin\frac{\hbar}{2}\right)$, plugging in Equation 5.9, and summing over $\beta\in H_{2}^{+}(\mathbb{P},\mathbb{Z})$ in Equation 5.11, we have the following equality,
@@
(5.12) $\sum_{g\geq 0}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}Q^{\beta}=\left(2\sin\frac{\hbar}{2}\right)\sum_{g\geq 0}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}$
Therefore, it suffices to show that Equation 5.12 holds. Specializing to winding-1 open invariants, the first statement of Theorem 5.3 gives the equality,
$Z_{\widehat{\mathbb{P}}}|_{\pi^{*}\beta-C}=Z_{\mathbb{P}}(V)|_{\square}Z_{\mathbb{C}^{3}}(V)|_{\square}$
This is equivalently expressed in Equation 5.8 as,
$Q^{-C}\sum_{\lambda_{1},\ldots,\lambda_{k}}(-1)^{\sum_{i}|\lambda_{i}|}q^{\sum_{i}\kappa_{\lambda_{i}}}\prod_{i=1}^{k}Q^{|\lambda_{i}|\ell_{i}}\prod_{j=1}^{k+1}C_{\lambda^{T}_{j}\lambda_{(j+1)\text{mod }k+1}\varnothing}=Q^{E}Z_{\mathbb{P}}(V)|_{\square}Z_{\mathbb{C}^{3}}(V)|_{\square}$
Since the curve class $\pi^{*}\beta-C$ is primitive, we obtain the winding-1 invariants directly from the partition functions. Hence we have,
$Q^{-C}\sum_{g,\beta}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}=Q^{E}\left(\sum_{g,\beta}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}Q^{\beta}\right)Z_{\mathbb{C}^{3}}(V)|_{\square}$
After canceling $Q^{E}$ with $Q^{-C}$ by flop invariance [KM] and noticing that $Z_{\mathbb{C}^{3}}(V)|_{\square}:=C_{\square\varnothing\varnothing}=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}$ (from pg.53, [AKMV]), we have
$\sum_{g,\beta}O_{g}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))\hbar^{2g-1}Q^{\beta}=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})\sum_{g,\beta}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}$
The result follows after making the variable change $q=e^{i\hbar}$, and redefining the generating function $\sum_{g,\beta}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)\hbar^{2g-2}Q^{\beta}$ up to a global factor of $i$. ∎
@@
In genus-0, we have $O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))=-n_{0}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ and $N_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=n_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ since $\pi^{*}\beta-C$ is a primitive curve class. By remarks made in Section 2.2 of [GRZ], $O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ equals the genus-0, open Gromov-Witten invariant of a moment fiber of $K_{\mathbb{P}}$. Thus, Corollary 5.4 is equivalent to Theorem 1.1, [LLW] in genus-0.
@@
For toric Fano manifolds, the genus-0, open Gromov-Witten invariant of a moment fiber was defined in the symplectic category by [FOOO]. As noted in Remark 5.5, both the open invariant of an outer AV-brane $L\subset K_{\mathbb{P}}$ or of a moment fiber agree in genus-0. By Corollary 5.4 and Theorem 1.1, [LLW], genus-0 open invariants defined either using stable relative maps as in [LLLZ] [FL] or in the symplectic category agree since they are both equal to $N_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$.

### v2

@@ 5.3. Open/closed correspondence
By flop invariance of the Topological Vertex [KM], we have $Q^{-C}=Q^{E}$, and the equality of closed partition functions,
$Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})=Z_{W}(Q^{\ell_{i}},Q^{E})$
The gluing or degeneration formula in the Topological Vertex ([LLLZ], Proposition 5.8) gives,
$Z_{W}(Q^{\ell_{i}},Q^{E})$ $=\sum_{\nu}(-1)^{|\nu|}Z_{\mathbb{P}}|_{\nu}C_{\nu^{T}\varnothing\varnothing}Q^{|\nu|E}$
$=\sum_{\nu}(-1)^{|\nu|}Z_{\mathbb{P}}Z_{\nu}C_{\nu^{T}\varnothing\varnothing}Q^{|\nu|E}$
which implies,
(5.9) $\frac{Z_{\widehat{\mathbb{P}}}(Q^{\ell_{i}},Q^{-C})}{Z_{\mathbb{P}}(Q^{\ell_{i}})}=\sum_{\nu}(-1)^{|\nu|}Z_{\nu}C_{\nu^{T}\varnothing\varnothing}Q^{|\nu|E}=1+\sum_{\nu\neq 0}(-1)^{|\nu|}Z_{\nu}C_{\nu^{T}\varnothing\varnothing}Q^{|\nu|E}$
Taking the log of Equation 5.9 to obtain connected invariants, we obtain the first expression. Since we divide by $Z_{\mathbb{P}}$, we have $w>0$ in the left generating series in Theorem 5.3. The second expression of the claim follows from a combinatorial analysis. ∎
We provide some examples of Theorem 5.3. For explicit values of LMOV invariants and Gopakumar-Vafa invariants for $\mathbb{P}=\mathbb{P}^{2}$, we refer to tables in Appendix A.
Example 5.4 (Theorem 5.3 in winding-1).
Setting $w=1$ in Theorem 5.3, we have,
$Q^{-C}\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)Q^{\pi^{*}\beta}\hbar^{2g-2}$ $=-Z_{(1)}C_{(1)\varnothing\varnothing}Q^{E}$
$=-(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-2}\widehat{f}_{(1)}Q^{E}$
The Gopakumar-Vafa formula [GV1] [GV2] for the curve class $\pi^{*}\beta-C$ is,
$\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)Q^{\pi^{*}\beta-C}\hbar^{2g-2}=\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}(-1)^{g-1}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)Q^{\pi^{*}\beta-C}(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2g-2}$
where $q=e^{i\hbar}$. Hence multiplying by $-(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$ and cancelling $Q^{-C}$ with $Q^{E}$, we have
(5.10) $\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}(-1)^{g}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)Q^{\pi^{*}\beta}(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2g}=\widehat{f}_{(1)}$
under the variable change $z=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}$. As the winding-1, open-BPS invariants are equal to the LMOV invariants in representation $\square$ by Equation 5.8, we have,
@@
For $w=1$, Example 5.4 of Theorem 5.3 implies,
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ $=(-1)^{g}N^{\mathrm{LMOV}}_{g,(1)}(K_{\mathbb{P}}/L,\beta)$
$=(-1)^{g}n_{g}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$
Example 5.6 (Theorem 5.3 in winding-2).
Setting $w=2$ in Theorem 5.3, we have
@@
$Q^{-2C}\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-2C)Q^{\pi^{*}\beta}\hbar^{2g-2}=\left(Z_{(2)}C_{(1,1)}+Z_{(1,1)}C_{(2)}-\frac{1}{2}(Z_{(1)}C_{(1)})^{2}\right)Q^{2E}$
The relevant quantities are described in [AKMV] and are,
$C_{(2)}$ $=\frac{q^{2}}{(q-1)(q^{2}-1)}$
$C_{(1,1)}$ $=\frac{q}{(q-1)(q^{2}-1)}$
$C_{(1)}$ $=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}$
$Z_{(2)}$ $=f_{(2)}(q,Q)+\frac{1}{2}f_{(1)}(q^{2},Q^{2})+\frac{1}{2}f_{(1)}(q,Q)^{2}$
$Z_{(1,1)}$ $=f_{(1,1)}(q,Q)-\frac{1}{2}f_{(1)}(q^{2},Q^{2})+\frac{1}{2}f_{(1)}(q,Q)^{2}$
$f_{(2)}$ $=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}(q^{\frac{-1}{2}}\widehat{f}_{(2)}-q^{\frac{1}{2}}\widehat{f}_{(1,1)})$
$f_{(1,1)}$ $=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}(-q^{\frac{1}{2}}\widehat{f}_{(2)}+q^{\frac{-1}{2}}\widehat{f}_{(1,1)})$
$Z_{(1)}$ $=f_{(1)}=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}\widehat{f}_{(1)}$
where $\widehat{f}$ are defined in Equation 5.7. We have $C_{(1,1)}+C_{(2)}=(q^{1/2}-q^{-1/2})^{-2},C_{(1,1)}-C_{(2)}=\frac{-1}{2(q-q^{-1})}$. The Gopakumar-Vafa formula for curve class $\pi^{*}\beta-2C$ is,
@@
(5.11) $\begin{split}Q^{\pi^{*}\beta-2C}\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}N_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-2C)\hbar^{2g-2}&=\sum_{\begin{subarray}{c}g\geq 0,\\ \beta\in NE(\mathbb{P})\end{subarray}}\left[(-1)^{g-1}n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-2C)\left(q^{\frac{1}{2}}-q^{\frac{-1}{2}}\right)^{2g-2}\right.\\ &+\left.(-1)^{g-1}n_{g}(K_{\widehat{\mathbb{P}}},\frac{\pi^{*}\beta}{2}-C)\frac{1}{2}(q-q^{-1})^{2g-2}\right]Q^{\pi^{*}\beta-2C}\end{split}$
Plugging in expressions, we have,
(5.12) $\begin{split}Z_{(2)}C_{(1,1)}+Z_{(1,1)}C_{(2)}-\frac{1}{2}(Z_{(1)}C_{(1)})^{2}&=f_{(2)}\frac{q}{(q-1)(q^{2}-1)}+f_{(1,1)}\frac{q^{2}}{(q-1)(q^{2}-1)}\\ &-\frac{1}{2(q-q^{-1})}f_{(1)}(q^{2},Q^{2})\\ &=(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}\left(q^{\frac{-1}{2}}\widehat{f}_{(2)}-q^{\frac{1}{2}}\widehat{f}_{(1,1)}\right)\frac{q}{(q-1)(q^{2}-1)}\\ &+(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{-1}\left(-q^{\frac{1}{2}}\widehat{f}_{(2)}+q^{\frac{-1}{2}}\widehat{f}_{(1,1)}\right)\frac{q^{2}}{(q-1)(q^{2}-1)}\\ &-\frac{1}{2(q-q^{-1})^{2}}\widehat{f}_{(1)}(q^{2},Q^{2})\\ &=\frac{-1}{(q^{\frac{1}{2}}-q^{\frac{-1}{2}})^{2}}\widehat{f}_{(2)}(q,Q)-\frac{1}{2(q-q^{-1})^{2}}\widehat{f}_{(1)}(q^{2},Q^{2})\end{split}$
Setting Equation 5.11 equal to Equation 5.12 implies, on the level of individual invariants,
Corollary 5.7.
For $w=2$, Example 5.6 of Theorem 5.3 implies,
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-2C)$ $=(-1)^{g}N^{\mathrm{LMOV}}_{g,(2)}(K_{\mathbb{P}}/L,\beta)$
For $\mathbb{P}=\mathbb{P}^{2},\widehat{\mathbb{P}}=\mathbb{F}_{1}$, Corollaries 5.5, 5.7 agree with values of Gopakumar-Vafa invariants of $K_{\mathbb{F}_{1}}$ in [HKR], Appendix B and LMOV invariants of $L\subset K_{\mathbb{P}^{2}}$ in [AKMV], §9.4. We conjecture that,
Conjecture 5.8.
For $w\in\mathbb{Z}_{>0}$, $\beta\in NE(\mathbb{P})$ and all $g\geq 0$,
$n_{g}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-wC)=(-1)^{g}N^{\mathrm{LMOV}}_{g,(w)}(K_{\mathbb{P}}/L,\beta)$
with proof in $w=1,2$.
Proof.
Refer to Examples 5.4, 5.6. ∎
Example 5.9.
In Appendix A, we provide tables of LMOV and Gopakumar-Vafa invariants that we computed to verify Conjecture 5.8 for $\mathbb{P}=\mathbb{P}^{2}$ and $w\leq 4,d\leq 5$ and $g\leq 3$. For example, the genus-$3$ Gopakumar-Vafa invariant of $K_{\mathbb{F}_{1}}$ in curve class $\pi^{*}5H-2C$ is $-992$ as listed in Table 1(b), and the genus-3, degree-5 LMOV invariant of $L\subset K_{\mathbb{P}}$ in representation $(2)$ is $992$.
@@
In genus-0, we have $O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))=n_{0}^{open}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ (see footnote in Example 5.14) and $N_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)=n_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$ since $\pi^{*}\beta-C$ is a primitive curve class. By remarks made in [GRZ], Section 2.2, $O_{0}(K_{\mathbb{P}}/L,\beta+\beta_{0},(1))$ equals the genus-0, open Gromov-Witten invariant of a moment fiber of $K_{\mathbb{P}}$. Thus, Corollary 5.5 is equivalent to [LLW], Theorem 1.1 in genus-0.
@@
For toric manifolds (both Fano and non-Fano), the genus-0, open Gromov-Witten invariant of a moment fiber was defined using symplectic geometry by [FOOO]. Both the open invariant of an outer AV-brane $L\subset K_{\mathbb{P}}$ and of a moment fiber in $K_{\mathbb{P}}$ agree in genus-0, by remarks in [GRZ], Section 2.2. Hence by Corollary 5.5 and [LLW], Theorem 1.1, genus-0 open Gromov-Witten invariants defined with either stable relative maps in algebraic geometry [LLLZ] [FL] or with symplectic geometry [FOOO] agree, as they both equal $N_{0}(K_{\widehat{\mathbb{P}}},\pi^{*}\beta-C)$.